### v1

@@ 3.2 Comparative Analysis of Pulse Characteristics
Risetime (R1): The arithmetic mean and standard deviation of the 10-90% risetime are s for Category A pulses, which aligns with values of s reported for Classical PBPs in a recent study by [21] on intracloud (IC) flashes;
Full width half max (FWHM): Our measured average of s also matches [21]’s s;
Zero-crossing Time (T1): The average first half-cycle duration of s, though seemingly smaller than the s measured by [21], actually aligns when their standard deviation is considered;
This statistical comparison to recent literature is presented in Table 4 (See Table 3 of [20] for more details). Similarly, Category B pulses share temporal characteristics with Narrow PBPs, with typical durations an order of magnitude shorter than Category A/Classical PBPs. [11], 2009 classified them as s, and our sub-microsecond measured risetimes and half-widths depicted in Table 2 are consistent with this.
@@
Duration (T1+T2) [s] | 26.2 | 21.6 22footnotemark: 2 | |
25 | 21 33footnotemark: 3 | 18.0 | 15.3 | |
17.2 | 11.9 44footnotemark: 4 | ||
@@
The observation of inverted polarity in approximately 10% of the bipolar Category A pulses (negative initial half-cycle)555These pulses were excluded from the primary statistical analysis due to their relative rarity. also parallels similar findings in downward lightning studies by [13], further supporting the connection between these phenomena.
Interestingly, we observed both Category A and B pulses occurring simultaneously during the early stage of leader development, followed by a later stage dominated exclusively by Category B pulses. This pattern parallels observations in downward negative flashes reported by [16], in which, as discussed in the previous section, they noted a bimodal distribution of the stepping process involving both long (200+ m) and short (10+ m) steps associated with the B and L phases, respectively (see their figures 1 & 6). The bimodal distribution of stepping processes observed in our data–with both Category A and B pulses present initially, followed by predominantly Category B pulses–mirrors the pattern described by [16] for downward negative flashes, as shown in Figure 7. This suggests that similar space leader formation and attachment processes may be occurring in both upward and downward leader development.
It is important to note, however, that the measured 2D step lengths for Classical PBPs are an order of magnitude larger than those measured for Category A pulses (25 m), while our estimated Category B step lengths (55 m) are of the same order as those reported during the L phase (Narrow PBPs), albeit a bit longer [16]. The former discrepancy is likely due to a difference in electric field geometry at the initiation point (in-cloud vs. tower tip), while the latter similarity is possibly due to the physical correspondence of downward stepping leaders approaching the ground (albeit of different polarities).
@@ 3.3 Physical Mechanism Comparison
As discussed in Section 1, there has been debate as to whether or not the discharge processes involved in the B and L phases of downward negative leaders are the same [3, 17]. Despite the clear similarities between Category A/B and Classical/Narrow PBPs, the degree to which the physical mechanisms at play are comparable is less evident. Nevertheless, some important common ground can be inferred.
@@
Aside from the initiation point differences (in-cloud vs. tower tip) mentioned in Section 3.2, the initial symmetry between upward and downward negative leaders readily explains the observed similarities between the pulses associated with their initial breakdown and propagation. These further suggest that the difference in conditions, specifically the availability of free charges in upward lightning due to the presence of a conducting ground and tower, as opposed to initiation within the thundercloud for downward lightning, may not play as important a role as one might expect, at least during the Breakdown phase.
@@
Since all observed E-field pulses are necessarily generated by an acceleration of charge, the lack of any corresponding current pulses measured at the tower base in Category B electric field pulses suggests that they are due to electrical activity disconnected from the stepping of the upward negative leader responsible for the generation of Category A pulses. Analysis of the HSC frames available for Flash #8 supports this interpretation. Figure 5 shows that the late stage Category B pulses starting at about 13.23 ms are generated by the stepping of a downward-connecting recoil leader prior to the main pulse. (See [15] for further discussion of how the Main Pulse of this Type 1 UPF (therein called UP2) appears to be triggered by the downward-stepping recoil leader connecting to the main current-carrying channel).888Note that Type 2 UPF Category B pulses are likely generated by a similar process that doesn’t connect to the main channel (or at least not in the same way). This is consistent with the observed domination of Category B and Narrow PB pulses at later times, when recoil/downward-connecting leaders become more common.
@@ 3.4 Validation of field–current Relationship Models for Classical PBPs
The simultaneous measurements of current and electric field waveforms provide a unique opportunity to validate existing models of the field–current relationship for PBPs. [7] proposed a physical model describing the PBP stage in downward negative lightning based on a predefined thundercloud charge structure, and deduced a proportionality between the peak currents and electric field amplitudes. As presented in Figure 6, we found a very strong linear correlation (, , ) between peak E-field and current amplitudes, with a best-fit .
After accounting for the known E-field enhancement factor of due to the topography between Mt. Säntis and Herisau [9], our measured peak E-field and current data of Category A pulses yielded V/m/kA, which is consistent with the field–current ratio of V/m/kA predicted by the model proposed in [7]. Our validation of this relationship using simultaneous measurements lends further support to the hypothesis of analogous physical processes for upward and downward negative leaders.
This field–current correlation also confirms suspicions of a different physical mechanism for Category B pulses, as their normalised E-field amplitudes of V/m would imply corresponding current pulse amplitudes ranging from 440 A [8] to 1.35 kA [7], depending on which field–current relation is used, which we do not observe (see Figure 5(b)).
4 Methods
The Mt. Säntis Lightning Research Facility, shown in Figure 7, is situated at 2502 m ASL in the Appenzell Alps of north-eastern Switzerland, and experiences 100 direct lightning strikes per year to its 124-meter-tall tower, which is equipped with a comprehensive current measurement system consisting of Rogowski coil and B-dot sensor pairs at two different heights: 24 and 82 meters above ground level (AGL). The system has a sampling rate of 50 MHz. Five kilometers away, atop Mt. Krönberg (1663 m ASL), is a high-speed camera (HSC) operating at 24,000 fps, with an exposure time of s and a resolution of 512512 pixels. Fifteen kilometres away, atop a building in Herisau, Switzerland (771 m ASL) with line-of-site, lies an flat-plate electric field antenna with a frequency range of 40 Hz to 40 MHz, a sampling rate of 10 MHz, and time constants of 4.2 ms (in 2014) and 8.4 ms (2021). This E-field signal is synchronized with the tower current signal by GPS timestamp if the antennae are functional at the time of the flash. If not, manual synchronization can be carried out via waveform matching, with an uncertainty of s. More detailed information on the Säntis measurement system can be found in [18].
@@ 5 Summary and conclusions
This study investigated the characteristics of stepped leader pulses associated with upward positive flashes at the Säntis tower, finding remarkable similarities with preliminary breakdown pulses observed in downward negative flashes. The key findings can be summarized as follows:
@@
The waveform characteristics and temporal behavior of Category A pulses associated with upward negative stepped leaders are very similar to “Classical” PBPs in downward lightning. This similarity suggests that the latter are generated during the 1st (breakdown) stage, i.e., creation of the downward negative leader;
@@
Category B pulses, whose lack of current correlation suggested an association with disconnected processes, were revealed by high-speed camera imaging to be the result of a downward-stepping recoil leader. This, in combination with their similarity to “Narrow” PBPs in downward lightning, suggests that the latter can be caused by recoil leader connections as well;
@@
We observed significant correlations between various bipolar Category A pulse parameters, including linear relationships between: (a) first and second E-field half-cycle durations (), (b) half-cycle peak amplitudes (), (c) E-field pulse peak and maximum derivative (also observed in Category B pulses), and (d) peak E-field and current amplitudes ();
@@
The theoretical relationship between E-field peak and current peak developed for PBPs in [7] was found to be consistent with our simultaneous measurements of currents and E-fields associated with Category A pulses in upward negative leaders;
@@
The temporal evolution of observed pulses showed a pattern similar to that observed in downward flashes, with both Category A and B pulses present initially, followed by predominantly Category B pulses, supporting the bimodal stepping process model proposed by [16].
These findings have important implications for lightning research. Despite physical differences in initiation environments (e.g., the presence of free charges in upward lightning from the conducting ground and tower, versus initiation within the thundercloud for downward lightning), the demonstrated similarities between upward and downward lightning initiation processes suggest common underlying physical mechanisms, especially in the case of high-altitude initiated upward flashes, where the breakdown voltage is similar, as in the case of the Säntis tower. Furthermore, our validation of the field–current relationship provides quantitative support for the [7] model of PBPs.
Perhaps most significantly, this study demonstrates that simultaneous current and electric field measurements, combined with high-speed camera observations of upward positive flashes can provide valuable insights into the initiation and development of downward negative flashes. The direct current measurements obtainable in tower-initiated upward leader pulses, when combined with electric field observations, offer a unique window into the still-debated lightning initiation processes that cannot be observed through conventional field measurements of downward lightning alone.
Future studies with more video and/or interferometric data, along with modelling efforts, are recommended to explore the distinct physical origins of these pulse types. Emphasis should be placed on expanding this dataset to include more events with complete measurement sets (particularly for Category B pulses), to further refine our understanding of the mechanisms underlying these processes and their implications for lightning initiation theory.
Acknowledgements
This work was supported in part by the Swiss National Science Foundation (Project no.s 200020_175594 and 200020_204235) and the European Union’s Horizon 2020 research and innovation program (grant agreement no. 737033-LLR). The authors would like to thank Florent Aviolat for developing a data-visualization software that expedited the identification of events, Kevin Abou Jaoude for his assistance with pulse analysis, and Hannes Kohlmann for the providing the EUCLID/ALDIS data for prior activity verification.
@@ Acknowledgements
All processed data analysed during this study are included in this published article (and its supplementary information files). Raw data sets generated during the current study are available from the corresponding author on reasonable request. (Link)
@@
Upon request from corresponding author. (Link)

### v2

@@ 4.2.2 Comparative Analysis of Pulse Characteristics
Risetime (R1): The arithmetic mean and standard deviation of the 10-90% risetime are s for Category A pulses, which aligns555We infer relative statistical “alignment” as each mean being within the other’s standard deviation. with values of s reported for Classical PBPs in a recent study by [shi_characterization_2024] on intracloud (IC) flashes;
Full width half max (FWHM): Our measured average of s also matches [shi_characterization_2024]’s s;
Zero-crossing Time (T1): The average first half-cycle duration of s, on par with [wu_preliminary_2013]’s s, though seemingly smaller than the s measured by [shi_characterization_2024], actually fits nicely within their standard deviation of s;
Pulse Duration (T1+T2): Bipolar Category A pulses have an average duration of s, falling within the range of 5 to 100+ s (with an average of s) reported for Classical PBPs in negative cloud-to-ground (CG) flashes [wang_beijing_2016, zhu_study_2016, granados_characterization_2022].
This statistical comparison to recent literature is presented in Table 4 (See Table 3 of [sekehravani_preliminary_2025] for more details). The inverted IC flashes of [shi_characterization_2024] have been included for their qualitative (and statistical) similarities, including “downward PB propagation”, a negative initial E-field change, and comparable temporal characteristics (as we show). Similarly, Category B pulses share temporal characteristics with Narrow PBPs, with typical durations an order of magnitude shorter than Category A/Classical PBPs. [nag_pulse_2008], 2009 classified them as s, and our sub-microsecond measured risetimes and half-widths depicted in Table 2 are consistent with this. A visual comparison of Category A/B and Classical/Narrow PB pulses is presented in Figure 5.
@@
| 6.1 | — 22footnotemark: 2 | ||
| Duration (T1+T2) [s] | 11 | — 33footnotemark: 3 | |
| 26.2 | 21.6 44footnotemark: 4 | 18.0 | 15.3 | |
| 25 | 21 55footnotemark: 5 | ||
| 17.2 | 11.9 66footnotemark: 6 | ||
@@
The observation of inverted polarity in approximately 10% of the bipolar Category A pulses (negative initial half-cycle)666These seven pulses were excluded from the primary statistical analysis due to their relative rarity. also parallels similar findings in downward lightning studies by [ogawa_initiation_1993], further supporting the connection between these phenomena.
Interestingly, we observed both Category A and B pulses occurring simultaneously during the early stage of leader development, followed by a later stage dominated exclusively by Category B pulses. This pattern parallels observations in downward negative flashes reported by [petersen_high-speed_2015], in which, as discussed in the previous section, they noted a bimodal distribution of the stepping process involving both long (200+ m) and short (10+ m) steps associated with the B and L phases, respectively (see their figures 1 & 6). The bimodal distribution of stepping processes observed in our data–with both Category A and B pulses present initially, followed by predominantly Category B pulses–mirrors the pattern described by [petersen_high-speed_2015] for downward negative flashes, as shown in Figure 8. This suggests that similar space leader formation and attachment processes may be occurring in both upward and downward leader development.
It is important to note, however, that the measured 2D step lengths for Classical PBPs are an order of magnitude larger than those measured for Category A pulses (25 m), while our estimated Category B step lengths (55 m) are of the same order as those reported during the L phase (Narrow PBPs), albeit a bit longer [petersen_high-speed_2015]. The former discrepancy is likely due to a difference in electric field geometry at the initiation point (in-cloud vs. tower tip), while the latter similarity is possibly due to the physical correspondence of downward stepping leaders approaching the ground (albeit of different polarities).
@@ 4.3 Physical Mechanism Comparison
As discussed in Section 1, there has been debate as to whether or not the discharge processes involved in the B and L phases of downward negative leaders are the same [clarence_preliminary_1957, proctor_vhf_1988]. Despite the clear similarities between Category A/B and Classical/Narrow PBPs, the degree to which the physical mechanisms at play are comparable is less evident. Nevertheless, some important common ground can be inferred.
@@
Aside from the initiation point differences (in-cloud vs. tower tip) mentioned in Section 4.2.2, the initial symmetry between upward and downward negative leaders readily explains the observed similarities between the pulses associated with their initial breakdown and propagation. These further suggest that the difference in conditions, specifically the availability of free charges in upward lightning due to the presence of a conducting ground and tower, as opposed to initiation within the thundercloud for downward lightning, may not play as important a role as one might expect, at least during the Breakdown phase. Note that the absence of the “initial E-change (IEC)” observed by [marshall_electromagnetic_2014] is to be expected here, as the IEC dampens with distance, such that it is effectively flat at 15 km.
@@
Since all observed E-field pulses are necessarily generated by an acceleration of charge, the lack of any corresponding current pulses measured at the tower base in Category B electric field pulses suggests that they are due to electrical activity disconnected from the stepping of the upward negative leader responsible for the generation of Category A pulses. Analysis of the HSC frames available for Flash #8 supports this interpretation. Figure 7 shows that the late stage Category B pulses starting at about 13.23 ms are generated by the stepping of a downward-connecting recoil leader prior to the main pulse. (See [oregelchaumont_underlying_2025] for further discussion of how the Type-1 Main Pulse of this particular UPF appears to be triggered by the downward-stepping recoil leader connecting to the main current-carrying channel.)999Note that Type 2 UPF Category B pulses are likely generated by a similar process that doesn’t connect to the main channel (or at least not in the same way). To the best of our knowledge, this is the clearest optical-electrical confirmation of a connection between the two phenomena to date, and is consistent with the observed domination of Category B and Narrow PB pulses at later times, when recoil/downward-connecting leaders become more common.
@@
It is important to remember, however, that only one flash recorded in this study had optical observations, and that the camera’s s exposure time, though relatively short, is still long enough that the 12 late-time Category B pulses were often grouped into single frames ( pulses per frame), as shown in Figure 7. It is for this reason that the approximate 2D step length of m was estimated by dividing the recoil leader’s measured channel length by the number of associated pulses, as described in Table 2. Furthermore, this is only one possible source of Category B pulses: those occurring at earlier times, especially when overlapping Category A pulses, are much more likely associated with the UNL stepping process, like narrow PBPs optically-confirmed to be generated by the stepping of downward negative leaders, as discussed in Section 4.2.2.
@@ 4.4 Validation of field–current Relationship Models for Classical PBPs
The simultaneous measurements of current and electric field waveforms provide a unique opportunity to validate existing models of the field–current relationship for PBPs. [kaspar_model_2017] proposed a physical model describing the PBP stage in downward negative lightning based on a predefined thundercloud charge structure, and deduced a proportionality between the peak currents and electric field amplitudes. As presented in Figure 8, we found a very strong linear correlation (, , ) between peak E-field and current amplitudes, with a best-fit .
After accounting for the known E-field enhancement factor of due to the topography between Mt. Säntis and Herisau [li_lightning_2016], our measured peak E-field and current data of Category A pulses yielded V/m/kA, which is consistent with the field–current ratio of V/m/kA predicted by the model proposed in [kaspar_model_2017]. Our validation of this relationship using simultaneous measurements lends further support to the hypothesis of analogous physical processes for upward and downward negative leaders.
This field–current correlation also confirms suspicions of a different physical mechanism for Category B pulses, as their normalised E-field amplitudes of V/m would imply corresponding current pulse amplitudes ranging from 440 A [kolmasova_subionospheric_2016] to 1.35 kA [kaspar_model_2017], depending on which field–current relation is used, which we do not observe (see Figure 7(b)).
@@ 5 Summary and conclusions
This study investigated stepped leader pulses associated with upward positive flashes at the Säntis tower and revealed strong similarities with preliminary breakdown pulses (PBPs) observed in downward negative flashes. The key findings can be summarised as follows:
@@
The waveform characteristics and temporal behavior of Category A pulses associated with upward negative stepped leaders are very similar to “Classical” PBPs in downward lightning, supporting the interpretation that the latter are generated during the initial breakdown stage associated with downward negative leader formation;
@@
Category B pulses, whose lack of current correlation suggests an association with disconnected processes, were revealed by high-speed camera imaging in one case to be the result of a downward-stepping recoil leader, which, in combination with their similarity to “Narrow” PBPs in downward lightning, indicates that at least a subset of the latter may also be produced by recoil leader connections;
@@
Significant correlations were observed among various bipolar Category A pulse parameters, including linear relationships between (a) first and second E-field half-cycle durations (), (b) half-cycle peak amplitudes (), (c) E-field peak and maximum derivative (also observed in Category B pulses), and (d) peak E-field and current amplitudes ();
@@
These simultaneous measurements of currents and E-fields associated with Category A pulses in upward negative leaders were found to be consistent with the peak field–current relationship proposed by [kaspar_model_2017], providing quantitative support for this theoretical model of PBPs;
@@
The temporal evolution of pulse types–initial coexistence of Categories A and B followed by a predominance of Category B–mirrors a pattern observed in downward flashes, supporting the bimodal stepping model proposed by [petersen_high-speed_2015].
These conclusions should be viewed in light of several limitations. The number of flashes with complete current, field, and high-speed video records is small, and the association of individual pulses with specific luminous features is constrained by camera spatio-temporal resolution. It is also important to acknowledge that in-cloud PBPs in downward flashes occur under different geometry, altitude, and ambient conditions than those near the tip of a grounded tower, which may influence breakdown thresholds, streamer inception, and leader stepping dynamics. Although the present findings are not meant to be interpreted as evidence of identical discharge physics under all atmospheric conditions, they nevertheless have important implications for lightning research: despite physical differences in initiation environments (e.g., the presence of free charges in upward lightning from the conducting ground and tower, versus initiation within the thundercloud for downward lightning), the demonstrated similarities between upward and downward negative leaders suggest common underlying physical mechanisms, especially in the case of high-altitude initiated upward flashes, where the breakdown voltage is similar, as in the case of the Säntis tower.
The availability of simultaneous channel-base current and remote electric field measurements, supplemented by high-speed imaging, makes tower-initiated upward flashes a potential proxy for investigating aspects of downward flash initiation that are otherwise inaccessible to conventional field measurements alone. Such measurements provide valuable constraints on still-debated models of lightning initiation and early leader development. Future studies incorporating larger and more complete datasets (particularly for Category B pulses), VHF interferometric observations, and/or 3D reconstructions are recommended to explore the distinct physical origins of these pulse types. Coupled with refined modelling of leader stepping and field–current relationships, these efforts will serve to further refine our understanding of the mechanisms underlying these processes and their implications for lightning initiation theory.
Acknowledgements The authors would like to thank Florent Aviolat for developing a data-visualisation software that expedited the identification of events, Kevin Abou Jaoude for his assistance with pulse analysis, and Hannes Kohlmann for the providing the EUCLID/ALDIS data for prior activity verification.
@@
Funding
This work was supported in part by the Swiss National Science Foundation (Project no.s 200020_175594 and 200020_204235) and the European Union’s Horizon 2020 research and innovation program (grant agreement no. 737033-LLR).
@@
All processed data analysed during this study are included in this published article. Raw data sets generated during the current study are available from the corresponding author on reasonable request.
@@
Upon request from corresponding author.